OpenStudy (anonymous):

Can you help me find the sum of series ((-1)^n) / (3n + 1) ?

OpenStudy (anonymous):

\[\sum_{n=0}^{Infinity}\frac{ (-1)^n }{ 3n +1 }\]

OpenStudy (amistre64):

have you tried writing out the first few partial sums to see if you can spot a pattern?

OpenStudy (anonymous):

Yes, I have, and I get a sequence \[1-\frac{ 1 }{ 4 } +\frac{ 1 }{ 7}-\frac{ 1 }{ 10 }+...+\frac{ (-1)^n }{ 3n+1 }\] which limit I can't figure out.

OpenStudy (amistre64):

the limit is not an elementary outcome .... what level of math is this?

OpenStudy (anonymous):

First year college.

OpenStudy (amistre64):

which course?

OpenStudy (anonymous):

Analysis 1, pure math

OpenStudy (amistre64):

hmmm the sum results in a transcendental number:\[\frac19(\sqrt{3}\pi+\log(8))\] so the best you could do is prolly approximating the limit. What are some methods at your disposal?

OpenStudy (anonymous):

I got slightly different solution using Wolfram Mathematica (there's also -9 in the brackets). Anyhow, I can use differentiation (though I'm not sure it'd help), MacLaurin polynomial/series and other basic methods.

OpenStudy (amistre64):

i believe differentation plays into the Maclaurin method .... finding successive derivatives to develop a polynomial representation. a stray thought i had was to use cos(n pi) instead of (-1)^n, find a relatively simply partial sum like S5 and average out the rest with integrations from 5 to inf, and 6 to inf

OpenStudy (amistre64):

S5 = 0.80728021978... (0.000829829-0.00116357)/2 = -0.0001668705 0.80728021978 -0.0001668705 = 0.80711334928 0.83564884826 but at S5, that doesnt seem to approximate it out too well

OpenStudy (anonymous):

This sum was needed to be calculated on the written part of the exam from a few years ago where we can't use calculators and such, so I think that there's a way to rewrite that fraction somehow in order for the sum to be easier to find.

OpenStudy (amistre64):

if we ignore the -1 and just write up a division \(1-3n+9n^2-27n^3+...\) ------------------ 1+3n ) 1 -(1+3n) -------- -3n -(-3n-9n^2) ---------- 9n^2 ..... but i cant see where that would be useful at the moment

OpenStudy (amistre64):

for this test, did you need an exact result?

OpenStudy (anonymous):

yes, unfortunately

OpenStudy (amistre64):

without ever have taken the course, im wondering if there is a relatively simple way to split out the positive and negative portions. but that end result is still daunting

OpenStudy (anonymous):

@oldrin.bataku could you look into this series, please?

OpenStudy (anonymous):

@amistre64 thank you for the effort :)

OpenStudy (amistre64):

thnx for the challenge :)

OpenStudy (amistre64):

some equivalent setups are \[\sum_0 \frac{1}{6n+1}-\frac{1}{6n+4}\] \[\sum_0 \frac{1}{6n+1}-\frac12\sum_0 \frac{1}{3n+2}\]

OpenStudy (anonymous):

wonder if splitting in to even and odd helps? probably not, just wondering

OpenStudy (anonymous):

then partial fractions or something?

OpenStudy (amistre64):

the final results are so obfuscated to me that a clear path is obscured to me

OpenStudy (anonymous):

yeah you got a \(\pi\) and log and a radical

OpenStudy (anonymous):

sounds like a party

OpenStudy (anonymous):

I've solved this :) The point was to use series \[\sum_{n=0}^{Infinity} \frac {(-1)^n} {3n+1} x ^{3n+1}\] and differentiate it to get \[\sum_{n=0}^{Infinity} (-1)^n x^{3n}\] which is MacLaurin's series for \[\frac{1} {1+x^3}\] Now find integral \[\int\limits_{0}^{x} \frac {1} {1+t^3}dt\] and it's going to be some function of x, and when we let \[\lim_{x \rightarrow 1}\] of that function it's going to give us the solution \[\frac{ 1 }{ 3 }\ln 2 + \frac {\pi} {3\sqrt3}\] P.S. I just had to share this with you :)

OpenStudy (amistre64):

interesting technique fer sure

OpenStudy (anonymous):

Right you should be able to recognize each term as a mere derivative. @amistre64 good idea but unfortunately the partial fractions will not telescope at all...